# BOTS: A Unified Framework for Bayesian Online Task Selection in LLM Reinforcement Finetuning

### v1

@@ 1 Introduction
Reinforcement finetuning (RFT) has become a key technique for aligning Large Language Models (LLMs) with human preferences and enhancing their reasoning capabilities (Jaech et al., 2024; Guo et al., 2025; Luo et al., 2025; Hu et al., 2025; Zeng et al., 2025). However, the effectiveness of RFT is highly sensitive to task selection (Parashar et al., 2025; Shen et al., 2025; Zhu et al., 2025; Wen et al., 2025; Li et al., 2025a). Naively training on a static, uniformly sampled dataset is inefficient: the model spends excessive computation on tasks that are either already mastered (too easy) or beyond reach (too hard) (Yu et al., 2025; Bae et al., 2025; Chen et al., 2025b). This inefficiency not only inflates training costs but also destabilizes optimization by reducing the effective batch size. The central challenge, therefore, is to dynamically select tasks of “just right” difficulty to maximize learning efficiency as the model’s capability evolves.
Existing methods to this challenge face several limitations. Offline task selection (Parashar et al., 2025; Shen et al., 2025; Zhu et al., 2025; Wen et al., 2025; Li et al., 2025a), which pre-schedules tasks from easy to hard, is too rigid and does not adapt to the evolving trajectory of the model. In response, a few online selection methods have been proposed, aiming to adaptively choose tasks based on model’s current capability. Core challenge of these methods lies in the tradeoff between the computational cost of collecting information and the accuracy of the resulting performance estimates. We argue that existing solutions are not sufficiently efficient: some expend excessive computation on information gathering, undermining efficiency, while others fail to fully exploit collected information, leading to suboptimal selection. On one hand, oversampling-based methods (Yu et al., 2025; Bae et al., 2025) find suitable tasks by rolling out oversized batches, introducing substantial extra cost. On the other hand, non-oversampling approaches typically rely on a single source of information—either leveraging historical evaluations as explicit evidence (Chen et al., 2025b; Qu et al., 2025) or exploiting inter-task correlations as implicit evidence (Sun et al., 2025). Our empirical results reveal a clear complementarity: explicit evidence provides stable and accurate task-difficulty estimates but suffers from a slow warm-up when historical evaluations are scarce in early training, whereas implicit evidence quickly guides early-stage selection yet becomes less reliable in later stages. These findings indicate that relying solely on one type of evidence leaves information underutilized and leads to suboptimal task selection. Therefore, a principled framework to fuse these complementary evidence sources is essential for robust and efficient online task selection.
@@ 2 Related Works
The impact of task difficulty on RFT of LLMs has become an active research topic. Inspired by the seminal idea of curriculum learning (Bengio et al., 2009), researchers have proposed various strategies for selecting appropriate tasks in LLM RFT. For instance, Parashar et al. (2025); Shen et al. (2025) advocate scheduling tasks from easy to hard, enabling LLMs to gradually acquire reasoning skills. Similar ideas have been extended to multi-modal LLMs (Zhu et al., 2025; Wen et al., 2025; Li et al., 2025a), though these methods mainly focus on offline task selection following an easy-to-hard trajectory.
More recently, online selection strategies have emerged, often targeting tasks of moderate difficulty. One line of work adopts sampling-based task filtering, where tasks with consistently trivial rewards (all zeros or ones) are considered uninformative and are down-weighted or filtered (Yu et al., 2025; Bae et al., 2025). While effective, these methods require additional rollouts, incurring non-trivial overhead. To avoid extra rollouts, several works attempt to predict task passing rates without direct rollouts. Chen et al. (2025b) formulate task selection as a non-stationary multi-armed bandit problem, treating each problem category (e.g., difficulty level or type) as an arm and using absolute advantage as a reward proxy, with posterior estimation based on historical outcomes. Qu et al. (2025) extend this framework to the task level. Although these methods eliminate extra rollouts, they rely solely on direct evaluations and overlook cross-task relationships. In contrast, Sun et al. (2025) propose evaluating a small set of reference tasks and predicting the passing rates of others using an attention-inspired kernel over embeddings. However, this approach still requires additional rollouts for the reference set and discards historical evaluation information.
@@ 3.3 Ultra-Light Interpolation Plug-in for Implicit Evidence
Computational Cost. The primary cost of our approach arises from evaluating the reference models for all tasks, which requires extensive rollouts. Fortunately, benchmarking standard base models has already become routine in RL dataset construction—for example, Cheng et al. (2025) report Qwen2.5-7B-Instruct (Yang et al., 2024) and Qwen3-30B-A3B (Yang et al., 2025) performance as meta-tags for filtering. Given these pre-computed references, the additional online overhead of our interpolation-based estimator is negligible compared to training-time rollouts (see Appendix F.1 for empirical results). This efficiency follows from two properties: (i) no extra rollouts are needed during training, and (ii) both updates and predictions reduce to lightweight vector operations.
@@ 3.4 Thompson Sampling for Task Selection
The first question is: at what difficulty level does the current model benefit most from training? Prior works (Chen et al., 2025b; Sun et al., 2025) show that, under binary rewards, tasks with success probability around $0.5$ are most informative for learning as they lead to gradients with larger expected magnitude than tasks with success probability close to 0 or $1$. We define the utility of a task as the absolute deviation of its posterior mean from a target success probability $p^{*}\in(0,1)$, with $p^{*}=0.5$ as the canonical choice.
Given this target success probability, the problem of online task selection naturally reduces to a non-stationary bandit problem. The central challenge is the tradeoff between exploitation and exploration: the model must decide whether to select tasks with high-confidence estimates close to the target rate in order to maximize immediate utility, or to select tasks with high uncertainty to gather information that may improve future decisions. A purely exploitative strategy might select tasks whose posterior mean $\tilde{p}_{t}^{k}$ is closest to $p^{*}$, but this risks overlooking tasks whose difficulty is currently uncertain but potentially optimal. To naturally balance exploration and exploitation, we employ Thompson Sampling (Thompson, 1933), a strategy renowned for both its empirical effectiveness and theoretical guarantees. More specifically, at each selection step $t$, we perform the following: (1) Posterior Sampling: Draw a sample of the passing rate from its current posterior distribution $\hat{p}_{k}\sim\mathrm{Beta}(\alpha_{t}^{k},\beta_{t}^{k})$ for each task $\mathcal{T}^{k}$ in the pool. (2) Selection: Select tasks with the highest estimated utilities $\{\hat{u}_{k}:=|\hat{p}_{k}-p^{*}|\}$ to form the training batch $\mathcal{B}_{t+1}$. This procedure elegantly prioritizes tasks that are likely to be near the target difficulty, while the inherent variance in sampling from the posterior ensures that tasks with higher uncertainty are naturally explored.
@@ 4.1 Setups
Dataset.  We conduct experiments on GURU (Cheng et al., 2025), a well-curated cross-domain RL dataset. Each subset is deduplicated, verified, and filtered. We use its math, code, and logic subsets (excluding the Zebra Puzzle due to its non-binary reward). Detailed information about the used datasets is provided in Appendix B.1 (for training) and Appendix B.2 (for evaluation).
RFT Setting.  We adopt GRPO (Shao et al., 2024), Qwen2.5-1.5B-Instruct and Qwen2.5-7B. Key hyperparameters include a learning rate of 1e-6, 16 rollouts per task, and a temperature of $1.0$. Comprehensive training details and used RL algorithm are provided in Appendix B.3 and B.5.
@@ 4.4 Performance Comparison Across Models and Domains
We conduct extended experiments across model sizes (1.5B and 7B) and task domains (math, code, logic). We compare BOTS under the default setting ($\lambda=0.1,\rho=0.1$, posterior sampling enabled) against four baselines. Two are out-of-framework baselines: (i) Random, where tasks are uniformly sampled, and (ii) Offline, where tasks are ranked from easy to hard based on success probabilities of Qwen2.5-7B-Instruct and Qwen3-30B-A3B (for tie-breaking). Two are within-framework baselines: (iii) BOTS-MoPPS, with $\lambda=0.0,\rho=0.0$ and posterior sampling enabled, which reduces our framework to MoPPS (Qu et al., 2025) and thus enables direct evaluation of a purely explicit-evidence strategy; and (iv) BOTS-DOTS, with $\lambda=1.0,\rho=1.0$ and posterior sampling disabled, serving as a proxy inspired by DOTS (Sun et al., 2025), which evaluates the long-term effectiveness of relying almost entirely on implicit evidence without corrective feedback. Implementation details and further discussion of these baselines are provided in Appendix B.6.
@@ B.1 Training Data
The training data for our experiments is sourced from GURU (Cheng et al., 2025), a well-curated, cross-domain RL dataset covering mathematics, code, logic, science, simulation, and tabular tasks. Each subset has been rigorously deduplicated, verified, and filtered. Our training utilizes the math, code, and logic subsets, with the Zebra Puzzle excluded from the logic portion due to its non-binary reward structure.
@@
The Code subset includes programming challenges from LeetCode (Xia et al., 2025), TACO-verified (Li, 2024), PrimeIntellect (Mattern et al., 2025), and LiveCodeBench (Jain et al., 2024). For PrimeIntellect and LiveCodeBench, it incorporates pre-filtered versions provided by DeepCoder [^2] https://www.together.ai/blog/deepcoder.
@@
The Logic subset is composed of several benchmarks designed to test structured reasoning. It includes Ordering Puzzles (relational ordering), Graph Puzzles (implicit graph traversal), the public training splits of ARC-AGI (Chollet et al., 2024) and ARC-AGI-2 (Chollet et al., 2025) (abstract grid transformations), and a 3.4k-sample from BARC (Li et al., 2024b) (synthetic ARC-style tasks). For these tasks, predictions are extracted from <answer> tags for reward calculation via exact match.
@@ B.2 Evaluation Benchmarks
Followed by GURU (Cheng et al., 2025), the evaluation suite consists of a set of established benchmarks to rigorously assess the model’s performance in a zero-shot setting across the same domains.
Math We evaluate mathematical reasoning on AIME24 (MAA, 2024) and MATH500 (Hendrycks et al., 2021), which cover a wide range of competition-level math problems.
Code Programming capabilities are assessed using HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and a subset of LiveCodeBench (Jain et al., 2024). These benchmarks span from basic function generation to complex algorithmic challenges.
Logic Logical and abstract reasoning are measured using Ordering Puzzles for general reasoning, and ARC-AGI (Chollet et al., 2024) for grid-based abstract reasoning.
@@ B.3 Training Details
The RFT setup follows standard practices (Cheng et al., 2025; Qu et al., 2025; Sun et al., 2025), employing GRPO (Shao et al., 2024) as the RL algorithm (detailed in Appendix B.5). All experiments are conducted with 8 NVIDIA A100 (80GB) GPUs. We utilize PyTorch’s FSDP for distributed training, with vLLM (Kwon et al., 2023) employed to accelerate the response generation during the rollout phase. The actor model is optimized using a learning rate of $1\times 10^{-6}$ and weight decay of 0.1. We apply a constant learning rate warmup for the first 10 steps and use gradient clipping with a maximum norm of 1.0. For the GRPO algorithm, we set the clipping ratio $\epsilon$ to 0.2 and generate 16 rollouts with a temperature of 1.0. The maximum prompt and response lengths are set to 4,096 and 8,192 tokens, respectively.
The training configurations are tailored for two LLMs: Qwen2.5-1.5B-Instruct and Qwen2.5-7B (Yang et al., 2024), across three reasoning domains: Math, Code, and Logic. The primary differences across settings lie in batch sizes and memory optimization strategies to accommodate different model sizes. Specifically, the 7B model experiments utilize a larger training batch size (512 vs. 256) and GRPO mini-batch size (64 vs. 32). To fit the larger model on the same hardware, we enable FSDP’s CPU offloading for all 7B model experiments. Key hyperparameters for all six experimental settings are summarized in Table 6.
@@ B.5 Policy Optimization Algorithm: GRPO
The policy is optimized using Group Relative Policy Optimization (GRPO) (Shao et al., 2024), a policy gradient method that operates without a learned value function. In GRPO, the advantage for a given response is computed by normalizing its reward against the statistics of a group of candidate responses sampled for the same prompt.
@@ B.6 Baseline Details
  • Setting from MoPPS (Qu et al., 2025): This ablation uses $\lambda=0.0,\rho=0.0$ with posterior sampling, relying solely on explicit evidence for task selection. Under this configuration, our framework reduces exactly to the setting studied in Qu et al. (2025), enabling a direct evaluation of a purely explicit-evidence-based strategy. Note that Qu et al. (2025) did not study the code or logic domains, nor did they provide principles for determining $\lambda$ ($1-\lambda$ in BOTS). We therefore reuse their hyperparameter choice for math ($\lambda=1$, i.e., $\lambda=0$ in BOTS) when evaluating other domains. This comparison is fair since no hyperparameter tuning is performed for either MoPPS or BOTS.
  • Proxy for DOTS (Sun et al., 2025): This ablation uses $\lambda=1.0,\rho=1.0$ with posterior sampling disabled, so task selection is driven entirely by implicit evidence from our estimator. It serves as a proxy for the approach in Sun et al. (2025), though our construction of pseudo-counts differs, see Appendix E.2 for more details. This baseline allows us to evaluate the long-term efficacy of a strategy that relies almost exclusively on implicit evidence without direct corrective feedback. We note that Sun et al. (2025) additionally explored reusing historical trajectories, which lies beyond the scope of this paper. To ensure clarity, we isolate such tricks from DOTS and focus solely on the online task selection component.

### v2

@@ 1 Introduction
Reinforcement finetuning (RFT) has become a key technique for aligning Large Language Models (LLMs) with human preferences and enhancing their reasoning capabilities (jaech2024openai; guo2025deepseek; luo2025deepscaler; hu2025open; zeng2025simplerl). However, the effectiveness of RFT is highly sensitive to task selection (parashar2025curriculum; shen2025skywork; zhu2025eduflow; wen2025structured; li2025temple). Naively training on a static, uniformly sampled dataset is inefficient: the model spends excessive computation on tasks that are either already mastered (too easy) or beyond reach (too hard) (yu2025dapo; bae2025online; chen2025self). This inefficiency not only inflates training costs but also destabilizes optimization by reducing the effective batch size. The central challenge, therefore, is to dynamically select tasks of “just right” difficulty to maximize learning efficiency as the model’s capability evolves.
Existing methods to this challenge face several limitations. Offline task selection (parashar2025curriculum; shen2025skywork; zhu2025eduflow; wen2025structured; li2025temple), which pre-schedules tasks from easy to hard, is too rigid and does not adapt to the evolving trajectory of the model. In response, a few online selection methods have been proposed, aiming to adaptively choose tasks based on model’s current capability. Core challenge of these methods lies in the tradeoff between the computational cost of collecting information and the accuracy of the resulting performance estimates. We argue that existing solutions are not sufficiently efficient: some expend excessive computation on information gathering, undermining efficiency, while others fail to fully exploit collected information, leading to suboptimal selection. On one hand, oversampling-based methods (yu2025dapo; bae2025online) find suitable tasks by rolling out oversized batches, introducing substantial extra cost. On the other hand, non-oversampling approaches typically rely on a single source of information—either leveraging historical evaluations as explicit evidence (chen2025self; qu2025can) or exploiting inter-task correlations as implicit evidence (sun2025improving). Our empirical results reveal a clear complementarity: explicit evidence provides stable and accurate task-difficulty estimates but suffers from a slow warm-up when historical evaluations are scarce in early training, whereas implicit evidence quickly guides early-stage selection yet becomes less reliable in later stages. These findings indicate that relying solely on one type of evidence leaves information underutilized and leads to suboptimal task selection. Therefore, a principled framework to fuse these complementary evidence sources is essential for robust and efficient online task selection.
@@ 2 Related Works
The impact of task difficulty on RFT of LLMs has become an active research topic. Inspired by the seminal idea of curriculum learning (bengio2009curriculum), researchers have proposed various strategies for selecting appropriate tasks in LLM RFT. For instance, parashar2025curriculum; shen2025skywork advocate scheduling tasks from easy to hard, enabling LLMs to gradually acquire reasoning skills. Similar ideas have been extended to multi-modal LLMs (zhu2025eduflow; wen2025structured; li2025temple), though these methods mainly focus on offline task selection following an easy-to-hard trajectory.
More recently, online selection strategies have emerged, often targeting tasks of moderate difficulty. One line of work adopts sampling-based task filtering, where tasks with consistently trivial rewards (all zeros or ones) are considered uninformative and are down-weighted or filtered (yu2025dapo; bae2025online). While effective, these methods require additional rollouts, incurring non-trivial overhead. To avoid extra rollouts, several works attempt to predict task passing rates without direct rollouts. chen2025self formulate task selection as a non-stationary multi-armed bandit problem, treating each problem category (e.g., difficulty level or type) as an arm and using absolute advantage as a reward proxy, with posterior estimation based on historical outcomes. qu2025can extend this framework to the task level. Although these methods eliminate extra rollouts, they rely solely on direct evaluations and overlook cross-task relationships. In contrast, sun2025improving propose evaluating a small set of reference tasks and predicting the passing rates of others using an attention-inspired kernel over embeddings. However, this approach still requires additional rollouts for the reference set and discards historical evaluation information.
@@ 3.3 Ultra-Light Interpolation Plug-in for Implicit Evidence
Computational Cost. The primary cost of our approach arises from evaluating the reference models for all tasks, which requires extensive rollouts. Fortunately, benchmarking standard base models has already become routine in RL dataset construction—for example, cheng2025revisiting report Qwen2.5-7B-Instruct (Yang2024Qwen25TR) and Qwen3-30B-A3B (yang2025qwen3) performance as meta-tags for filtering. Given these pre-computed references, the additional online overhead of our interpolation-based estimator is negligible compared to training-time rollouts (see Appendix F.1 for empirical results). This efficiency follows from two properties: (i) no extra rollouts are needed during training, and (ii) both updates and predictions reduce to lightweight vector operations.
@@ 3.4 Thompson Sampling for Task Selection
The first question is: at what difficulty level does the current model benefit most from training? Prior works (chen2025self; sun2025improving) show that, under binary rewards, tasks with success probability around $0.5$ are most informative for learning as they lead to gradients with larger expected magnitude than tasks with success probability close to 0 or $1$. We define the utility of a task as the absolute deviation of its posterior mean from a target success probability $p^{*}\in(0,1)$, with $p^{*}=0.5$ as the canonical choice.
Given this target success probability, the problem of online task selection naturally reduces to a non-stationary bandit problem. The central challenge is the tradeoff between exploitation and exploration: the model must decide whether to select tasks with high-confidence estimates close to the target rate in order to maximize immediate utility, or to select tasks with high uncertainty to gather information that may improve future decisions. A purely exploitative strategy might select tasks whose posterior mean $\tilde{p}_{t}^{k}$ is closest to $p^{*}$, but this risks overlooking tasks whose difficulty is currently uncertain but potentially optimal. To naturally balance exploration and exploitation, we employ Thompson Sampling (thompson1933likelihood), a strategy renowned for both its empirical effectiveness and theoretical guarantees. More specifically, at each selection step $t$, we perform the following: (1) Posterior Sampling: Draw a sample of the passing rate from its current posterior distribution $\hat{p}_{k}\sim\mathrm{Beta}(\alpha_{t}^{k},\beta_{t}^{k})$ for each task $\mathcal{T}^{k}$ in the pool. (2) Selection: Select tasks with the highest estimated utilities $\{\hat{u}_{k}:=|\hat{p}_{k}-p^{*}|\}$ to form the training batch $\mathcal{B}_{t+1}$. This procedure elegantly prioritizes tasks that are likely to be near the target difficulty, while the inherent variance in sampling from the posterior ensures that tasks with higher uncertainty are naturally explored.
@@ 4.1 Setups
Dataset.  We conduct experiments on GURU (cheng2025revisiting), a well-curated cross-domain RL dataset. Each subset is deduplicated, verified, and filtered. We use its math, code, and logic subsets (excluding the Zebra Puzzle due to its non-binary reward). Detailed information about the used datasets is provided in Appendix B.1 (for training) and Appendix B.2 (for evaluation).
RFT Setting.  We adopt GRPO (grpo2024), Qwen2.5-1.5B-Instruct and Qwen2.5-7B. Key hyperparameters include a learning rate of 1e-6, 16 rollouts per task, and a temperature of $1.0$. Comprehensive training details and used RL algorithm are provided in Appendix B.3 and B.5.
@@ 4.4 Performance Comparison Across Models and Domains
We conduct extended experiments across model sizes (1.5B and 7B) and task domains (math, code, logic). We compare BOTS under the default setting ($\lambda=0.1,\rho=0.1$, posterior sampling enabled) against four baselines. Two are out-of-framework baselines: (i) Random, where tasks are uniformly sampled, and (ii) Offline, where tasks are ranked from easy to hard based on success probabilities of Qwen2.5-7B-Instruct and Qwen3-30B-A3B (for tie-breaking). Two are within-framework baselines: (iii) BOTS-MoPPS, with $\lambda=0.0,\rho=0.0$ and posterior sampling enabled, which reduces our framework to MoPPS (qu2025can) and thus enables direct evaluation of a purely explicit-evidence strategy; and (iv) BOTS-DOTS, with $\lambda=1.0,\rho=1.0$ and posterior sampling disabled, serving as a proxy inspired by DOTS (sun2025improving), which evaluates the long-term effectiveness of relying almost entirely on implicit evidence without corrective feedback. Implementation details and further discussion of these baselines are provided in Appendix B.6.
@@ B.1 Training Data
The training data for our experiments is sourced from GURU (cheng2025revisiting), a well-curated, cross-domain RL dataset covering mathematics, code, logic, science, simulation, and tabular tasks. Each subset has been rigorously deduplicated, verified, and filtered. Our training utilizes the math, code, and logic subsets, with the Zebra Puzzle excluded from the logic portion due to its non-binary reward structure.
@@
The Code subset includes programming challenges from LeetCode (xia2025leetcodedataset), TACO-verified (likaixin2024taco-verified), PrimeIntellect (2025synthetic1), and LiveCodeBench (jain2024livecodebench). For PrimeIntellect and LiveCodeBench, it incorporates pre-filtered versions provided by DeepCoder [^2] https://www.together.ai/blog/deepcoder.
@@
The Logic subset is composed of several benchmarks designed to test structured reasoning. It includes Ordering Puzzles (relational ordering), Graph Puzzles (implicit graph traversal), the public training splits of ARC-AGI (chollet2024arc) and ARC-AGI-2 (chollet2025arc) (abstract grid transformations), and a 3.4k-sample from BARC (li2024barc) (synthetic ARC-style tasks). For these tasks, predictions are extracted from <answer> tags for reward calculation via exact match.
@@ B.2 Evaluation Benchmarks
Followed by GURU (cheng2025revisiting), the evaluation suite consists of a set of established benchmarks to rigorously assess the model’s performance in a zero-shot setting across the same domains.
Math We evaluate mathematical reasoning on AIME24 (aime2024) and MATH500 (hendrycks2021measuringmathematicalproblemsolving), which cover a wide range of competition-level math problems.
Code Programming capabilities are assessed using HumanEval (chen202humaneval), MBPP (austin2021mbpp), and a subset of LiveCodeBench (jain2024livecodebench). These benchmarks span from basic function generation to complex algorithmic challenges.
Logic Logical and abstract reasoning are measured using Ordering Puzzles for general reasoning, and ARC-AGI (chollet2024arc) for grid-based abstract reasoning.
@@ B.3 Training Details
The RFT setup follows standard practices (cheng2025revisiting; qu2025can; sun2025improving), employing GRPO (grpo2024) as the RL algorithm (detailed in Appendix B.5). All experiments are conducted with 8 NVIDIA A100 (80GB) GPUs. We utilize PyTorch’s FSDP for distributed training, with vLLM (kwon2023efficient) employed to accelerate the response generation during the rollout phase. The actor model is optimized using a learning rate of $1\times 10^{-6}$ and weight decay of 0.1. We apply a constant learning rate warmup for the first 10 steps and use gradient clipping with a maximum norm of 1.0. For the GRPO algorithm, we set the clipping ratio $\epsilon$ to 0.2 and generate 16 rollouts with a temperature of 1.0. The maximum prompt and response lengths are set to 4,096 and 8,192 tokens, respectively.
The training configurations are tailored for two LLMs: Qwen2.5-1.5B-Instruct and Qwen2.5-7B (Yang2024Qwen25TR), across three reasoning domains: Math, Code, and Logic. The primary differences across settings lie in batch sizes and memory optimization strategies to accommodate different model sizes. Specifically, the 7B model experiments utilize a larger training batch size (512 vs. 256) and GRPO mini-batch size (64 vs. 32). To fit the larger model on the same hardware, we enable FSDP’s CPU offloading for all 7B model experiments. Key hyperparameters for all six experimental settings are summarized in Table 6.
@@ B.5 Policy Optimization Algorithm: GRPO
The policy is optimized using Group Relative Policy Optimization (GRPO) (grpo2024), a policy gradient method that operates without a learned value function. In GRPO, the advantage for a given response is computed by normalizing its reward against the statistics of a group of candidate responses sampled for the same prompt.
@@ B.6 Baseline Details
  • Setting from MoPPS (qu2025can): This ablation uses $\lambda=0.0,\rho=0.0$ with posterior sampling, relying solely on explicit evidence for task selection. Under this configuration, our framework reduces exactly to the setting studied in qu2025can, enabling a direct evaluation of a purely explicit-evidence-based strategy. Note that qu2025can did not study the code or logic domains, nor did they provide principles for determining $\lambda$ ($1-\lambda$ in BOTS). We therefore reuse their hyperparameter choice for math ($\lambda=1$, i.e., $\lambda=0$ in BOTS) when evaluating other domains. This comparison is fair since no hyperparameter tuning is performed for either MoPPS or BOTS.
  • Proxy for DOTS (sun2025improving): This ablation uses $\lambda=1.0,\rho=1.0$ with posterior sampling disabled, so task selection is driven entirely by implicit evidence from our estimator. It serves as a proxy for the approach in sun2025improving, though our construction of pseudo-counts differs, see Appendix E.2 for more details. This baseline allows us to evaluate the long-term efficacy of a strategy that relies almost exclusively on implicit evidence without direct corrective feedback. We note that sun2025improving additionally explored reusing historical trajectories, which lies beyond the scope of this paper. To ensure clarity, we isolate such tricks from DOTS and focus solely on the online task selection component.